HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

SELF AND N_{2} COLLISIONAL BROADENING OF FAR-INFRARED METHANE LINES AT LOW TEMPERATURE WITH APPLICATION TO

TITAN

\underline{C. RICHARD} , V. BOUDON, Laboratoire Interdisciplinaire Carnot de Bourgogne, UMR 6303 CNRS - Université Bourgogne Franche-Comté, 9 Av. A. Savary, BP 47870, F-21078 Dijon Cedex, France; L. MANCERON, AILES, Synchrotron SOLEIL, L’Orme des Merisiers Saint-Aubin, F-91192 Gif-sur-Yvette, France; J. VANDER AUWERA, SQUARES, C.P. 160/09, Université Libre de Bruxelles,

B-1050 Brussels, Belgium; S. VINATIER, B. BÉZARD, M. HOUELLE, LESIA, Observatoire de Paris, Université PSL, Sorbonne Université, Université Paris Cité,

CNRS, 5 place Jules Janssen, 92195 Meudon, France

We report the measurement of broadening coefficients of pure rotational lines of methane at different pressure and temperature conditions. A total of 27 far-infrared spectra were recorded at the AILES beamline of the SOLEIL synchrotron at room-temperature, 200 K and 120 K, in a range of 10 to 800 mbar. Self and N_{2} broadening coefficients and temperature dependence exponents of methane pure rotational lines have been measured in the 73–136 cm^{−1} spectral range using multi-spectrum non-linear least squares fitting of Voigt profiles. These coefficients were used to model spectra of Titan that were compared to a selection of equatorial Cassini/CIRS spectra, showing a good agreement for a stratospheric methane mole fraction of (1.17 ± 0.08)%.

p-number: p053  Submitted on Fri, 16 Jun 2023 15:50:58 +0200